OpenStudy (anonymous):

The final exam in a sociology course consists of 100 multiple-choice questions. Each question has five possible answers, and only one of them is correct. An unprepared student makes random guesses for all the answers. Find the mean and standard deviations for the number of correct answers for this student.

OpenStudy (campbell_st):

ask @asnaseer, very good at probability and gives lots of help

Directrix (directrix):

In a Binomial Probabilty Distribution, the mean can be found by M = n * p where n is the number of trials ( 100 questions here) and p is the probability of success (1/5 here). So, M = 100 * 1/5 = 20

Directrix (directrix):

Standard Dev = √( (n*p*(1-p) ) where (1-p) is the probability of an incorrect answer Standard Dev = √ (100*(1/5)*(4/5) = √ 16 = 4 @mjasma Check this and see what you think.